Biotech Company is selling their product in single units for $4.17 each and in bulk bundles of 10 which cost $25.85 per bundle.
They sell 829 units of their product in one day which gives them a gross revenue of $2,426.68.
How many single units and how many bundles did they sell that day?
here is what i have so far:
4.17x + 25.85y = 2426.68
x + y = 829
for some reason i can't get the answer to work out.

Answer by doukungfoo(195)

your first equation is correct but the second one is wrong
It should be
x + 10y = 829
Reason is because y is equal to bundles of 10.
So to find out how many single units were sold in all the bundles sold
you have to multiply number of bundles (y) times 10 (# of single units in each bundle)
y = number of bundles (bundles contain 10 units)
x = single unit
Ok so here are the two equations:
4.17x + 25.85y = 2426.68
and
x + 10y = 829
Start with the second equation. Set it equal to x
x = 829 - 10y
now use this value of x in the first equation so
4.17(x) + 25.85y = 2426.68
4.17(829 - 10y) + 25.85y = 2426.68
3456.93 - 41.7y + 25.85y = 2426.68
3456.93 - 15.85y = 2426.68
-15.85y = 2426.68 - 3456.93
-15.85y = -1030.25
y = 65
Ok at this point we have solved for y which is the number of bundles
So 65 bundles
Now we can take this value for y and plug it into the second equation to
find the number of single units sold (x)
x + 10y = 829
x + 10(65) = 829
x + 650 = 829
x = 179
So the number of single units sold is 179


I got 65 bundles and 179 single units sold.
so
179 + 10(65) = 829